OpenStudy (adamaero):

4/(2+x) the same as -4/(x-2) ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

\[-\frac{ 4 }{ x-2 }=\frac{ -4 }{ -\left( -x+2 \right) }=\frac{ 4 }{ 2-x }\]

OpenStudy (anonymous):

4/(2+x) multiplied by -1 is -4/-(2+x) and if distribute the negative in the denominator properly, you get -4/(-2-x).